OpenStudy (anonymous):

Use the Event Root Property: (2x+3)² + 10 =110

mathslover (mathslover):

\[\large{(2x+3)^2=110-10}\]

Parth (parthkohli):

Well. Subtract 10 from both sides\[(2x + 3)^2 = 100 \]Find the square root of both sides.\[2x + 3 = \pm10 \]

Parth (parthkohli):

Solve these two equations:\[2x + 3 = +10 \]\[2x + 3 = -10 \]

Parth (parthkohli):

I believe that you'd manage the rest, no?

OpenStudy (anonymous):

yes thank you :)

Parth (parthkohli):

Heh. You're welcome!\[\Huge \ddot \smile \]

hero (hero):

You spelled "Event" wrong.

Parth (parthkohli):

I still don't understand what the 'event root property' is. I feel that it's the one that I explained, right?

hero (hero):

I think he meant "Even" Root Property

Parth (parthkohli):

What does it say? The same that I just posted?

hero (hero):

Yes, but in more general terms of course.

Parth (parthkohli):

Okay :)

hero (hero):

It's even root property because obviously, it's different from doing odd roots.

Parth (parthkohli):

What's that property?

hero (hero):

There's a distinct method for each root

hero (hero):

Just taking the n root of both sides, where n = odd number

Parth (parthkohli):

But do they still get the same solutions, or different? Merely two different ways to solve an equations, eh?

hero (hero):

They get the "correct" solutions because those are the required methods necessary to solve these kinds of equations.

Parth (parthkohli):

Hmm.

hero (hero):

It's nothing complicated or anything you don't already know.

hero (hero):

You were doing the methods and probably just didn't know the specific name for it.

Parth (parthkohli):

Yeah. Whatever they name these kinds of properties. Lol.

Parth (parthkohli):

Exactly.

OpenStudy (anonymous):

it was the even root*

hero (hero):

You can still change it. Click "Edit Question" button in the blue box above.